A first order rkn at 298K can be expressed  by generalisation logK(ln s-1)=17.64-(2.5*104/T) then what is the activation energy of the rkn will be [K=rate constant of rkn ;T=Kelvin temp.] ?

Grade:12

1 Answers

SPARSH JAUHARI
12 years ago

K=Ae^(-Ea/RT)
log k= logA -(Ea)/RT
here logA=17.64
and Ea/R=2.5*10^4
Put R=8.314